# Contracting with a Mechanism Designer

### v1

@@ 5 Impact of Intermediary and Price of Double Marginalization
To this end, we introduce the concept of the price of double marginalization (PoDM), which captures the efficiency loss resulting from the intermediary’s presence. Specifically, PoDM compares the principal’s utility under the second-best benchmark—where the intermediary optimizes on her behalf—to the utility achieved in our delegated mechanism design model.
@@
To evaluate the overall efficiency loss relative to the welfare-optimal outcome, we additionally consider the classical price of anarchy (PoA). This benchmark compares the principal’s utility under the first-best benchmark—which maximizes total welfare—to that in our model. The two benchmarks correspond, respectively, to the maximum welfare and the maximum revenue in a related single-item auction, as characterized in our earlier analysis.
@@
where $U^{*}$ denotes the principal’s optimal utility in our model, and $U_{\textsc{fb}}$ denotes the first-best benchmark, corresponding to the maximum expected social welfare achievable when the principal can directly select the agent that maximizes welfare:
@@
Recall that Lemma 2.1 shows that the virtual welfare maximizer mechanism ensures truthfulness and achieves maximum expected revenue. Since the contribution distributions $\bm{F}$ are regular by 2.4, we have the optimal mechanism for the principal is the virtual welfare maximizer mechanism. Therefore, the second-best benchmark is equivalent to the expected revenue of the virtual welfare maximizer mechanism, i.e., $U_{\textsc{sb}}=\mathbb{E}_{\bm{w}\sim\bm{F}}\!\left[{\max\nolimits_{i\in[n]}\left\{{\left(\phi_{i}(w_{i})\right)^{+}}\right\}}\right]$.
Furthermore, the first-best benchmark is equivalent to the maximized utility of the principal in the scenario where the principal has direct access to agents and knows their cost realizations. Indeed, the assumption in this scenario implies that the principal can assign the task to the one with the lowest cost, thereby obtaining a utility equivalent to her contribution. Therefore, the first-best benchmark is equivalent to the expectation of the maximum value between 00 and the first order statistic $w_{\max}$ from the $n$ contribution, i.e., $U_{\textsc{fb}}=\mathbb{E}_{\bm{w}\sim\bm{F}}\!\left[{{\left(w_{\max}\right)^{+}}}\right]$.
@@ 5.1 Comparison under Regular Distributions
In this subsection, we bound the two ratios under 5.4. Our main result in Theorem 5.1 shows that for a regular contribution distribution $F$, if the agents’ maximum possible contribution is at most $\kappa$ times their minimum possible contribution, then PoDM is tightly bounded by $\Theta(\log\kappa)$, while the PoA is tightly bounded by $\Theta((\log\kappa)^{2})$. [^14] We say a ratio is tight if, for all contribution distributions satisfying the given conditions and any number of agents $n$, the ratio does not exceed a certain value, while there exists at least one distribution and one $n$ where the ratio equals or approaches this value.
@@
Proof overview of Theorem 5.1 Here, we outline the main ideas behind the proof of Theorem 5.1, which consists of three parts. In the first part, we establish the upper bound for PoDM under two conditions with different arguments. For the case where $4\leq\bar{w}/\underline{w}\leq\kappa$, we apply the monotonicity of the virtual value to reduce the maximized utility $U^{*}$ (Lemma 5.3). We then divide the quantile space of any contribution distribution into $12\left\lceil\log\kappa\right\rceil+$ 1 pieces. Next, using the monotonicity of the revenue curve for regular distributions, we find an upper bound for monopoly revenue $R^{*}$ and the ratio between the maximized utility $U^{*}$ and $R^{*}$. For the case where $1\leq\bar{w}/\underline{w}\leq 4$, we construct a quantile $q^{{\dagger}}$ representing the largest quantile where $R^{\prime}(q^{{\dagger}})\geq\underline{w}/2$. Then, by rescaling the integral form for $R^{*}$, we obtain a constant upper bound for the ratio between $U^{*}$ and $R^{*}$. We finally use the ratio of revenues from anonymous pricing and Myerson’s optimal auction (Lemma 5.4) to establish an upper bound for the second-best benchmark $U_{\textsc{sb}}$. This process allows us to establish the upper bound for PoDM.
@@
In the second part, we define the truncated equal-revenue distribution (Definition 5.6) and show that it achieves the maximum first-best benchmark for a single agent (Lemma 5.5). Combining this with the upper bound for PoDM, we derive the upper bound for PoA.
@@
In the third part, we provide an example (Example 5.5) to illustrate the lower bounds of the two ratios and demonstrate that the two ratios are tight. Specifically, the lower bound part of Theorem 5.1 utilizes Example 5.5. The upper bound part utilizes Lemmas 5.2, 5.3, 5.4 and 5.5.
Example 5.5.
Lemma 5.2 (Feng and Jin, 2024, Theorems 1 and 2).
  1. 1.
  2. 2.
Lemma 5.3.
Proof.
We first observe that the virtual value function is weakly decreasing implies that the monopoly reserve price is weakly increasing. Indeed, if $\phi^{(n)}(z)\geq\phi^{(n+1)}(z)$ holds for any $z\in\mathtt{supp}(F)$, then by the definition of the reserve price, we have
Given that $F^{n}$ is regular, we also notice that $\phi^{(n)}(z)$ is non-decreasing with respect to $z$. This leads that
$\theta^{(n+1)}\geq\theta^{(n)}.$
$\phi^{(n)}(z)=z-\frac{1}{h^{(n)}}=z-\frac{1-F^{n}}{nF^{n-1}F^{\prime}}.$
$\phi^{(n)}(z)\geq\phi^{(n+1)}(z).$
This completes our proof. ∎
Lemma 5.4 (Yan, 2011, Theorem 4.1).
To sell a single item among multiple buyers with i.i.d. and regular distributions, the revenue of anonymous pricing is an $\eta$-approximation to that of Bayesian revenue-optimal mechanism, where $\eta=\frac{{\mathrm{e}}}{{\mathrm{e}}-1}$.
Definition 5.6 (The truncated equal-revenue distribution).
One can verify that the monopoly revenue of the truncated equal-revenue distribution $F^{{\dagger}}$ is also $R^{*}$ and $F^{{\dagger}}$ is also regular since the revenue curve of $F^{{\dagger}}$ is concave. Moreover, we notice that the CDF is discontinuous at $\bar{w}$ which corresponds to a point-mass of $R^{*}/\bar{w}$ at value $\bar{w}$.
Lemma 5.5.
Proof.
Let $F$ be an arbitrary regular distribution with support $\left[\underline{w},\bar{w}\right]$ and monopoly revenue $R^{*}\in[\underline{w},\bar{w}]$. Observe that the revenue curve of the truncated equal-revenue distribution $F^{{\dagger}}$ is $R^{{\dagger}}(q)=\min\left\{\bar{w}q,R^{*}\right\}$. Consider the revenue curve of distribution $F$, given by $R(q)=zq$, where $q=1-F(z)$. On the one hand, we know $R(q)\leq\bar{w}q$ since $z\in[\underline{w},\bar{w}]$; on the other hand, we have $R(q)\leq R^{*}$ by the definition of $R^{*}$. It follows that the revenue curve of distribution $F$ is always positioned below that of the truncated equal-revenue distribution $F^{{\dagger}}$, i.e., $R(q)\leq R^{{\dagger}}(q)$ for all $q\in[0,1]$. Consequently, since both $R(\cdot)$ and $R^{{\dagger}}(\cdot)$ are concave, for any $z\in[\underline{w},\bar{w}]$, we have
$1-F(z)\leq 1-F^{{\dagger}}(z).$
Now we are ready to prove Theorem 5.1.
Proof of Theorem 5.1..
We first show the upper bound parts utilizing the above lemmas, and then show the lower bound part by analyzing Example 5.5.
Upper bounding PoDM.  We first observe that $F^{n}$ is also regular by Lemma 5.2. Thus, we have $\phi^{(n)}(z)\geq\phi^{(n+1)}(z)$ for any $z$ in $\mathtt{supp}(F)$ by Lemma 5.3. Then we can obtain that
$U^{*}\geq\max\limits_{q\in[0,1]}q\cdot R^{\prime}(q).$ (2)
Let $K=2\left\lceil\log\kappa\right\rceil$. Divide the quantile space $[0,1]$ of $F^{n}$ into $K+1$ pieces by introducing the following points: $01>q_{0}>q_{1}>\cdots>q_{K}=$ 0. We set the slope of the revenue curve (the virtual value) $R^{\prime}(q_{0})=0$ and $R^{\prime}(q_{k})=\underline{w}\cdot 2^{k-2\left\lceil\log\kappa\right\rceil+\log\kappa}$ for every index $k\in[K]$. Note that $R^{\prime}(q_{K})=\bar{w}$. One can verify that this construction holds for any regular distribution $F^{n}$ since the virtual value function is non-decreasing. Consequently, for every $q\in[q_{k},q_{k-1}]$ with $k\in[2:K]$, it holds
and for every $q\in[q_{1},q_{0}]$, it holds
Since the support of $F$ is $[\underline{w},\bar{w}]$, we know $R(1)\geq\underline{w}$. Thus, we know that monopoly revenue $R^{*}=R(q_{0})\geq R(1)\geq\underline{w}$. Moreover, since $R(q_{0})=0$, we have
$R^{*}$ $={}\int_{0}^{q_{0}}R^{\prime}(q)\cdot{\mathrm{d}}q$
$={}\sum_{k\in[K]}\int_{q_{k}}^{q_{k-1}}R^{\prime}(q)\cdot{\mathrm{d}}q$
where inequality (a) holds by Equation 3 and inequality (b) holds by Equation 4. It follows that
which implies that there exists $k^{{\dagger}}\in[2:K]$ such that
Note that by Lemma 5.4, the second-best benchmark satisfies $R^{*}\leq U_{\textsc{sb}}\leq\eta\cdot R^{*}$. Combine Equation 2, we can obtain that
Finally, we consider the case where $1\leq\bar{w}/\underline{w}\leq 4$.
We let $q^{{\dagger}}$ be the largest quantile such that $R^{\prime}(q^{{\dagger}})\geq\underline{w}/2$ and $q^{*}$ be the monopoly quantile such that $R^{*}=R\left(q^{*}\right)$. Notice $q^{{\dagger}}$ is well-defined since $R^{\prime}(0)=\bar{w}$. By definition, we have $0\leq q^{{\dagger}}\leq q^{*}\leq 1$ and $R^{\prime}(q^{{\dagger}})\leq\underline{w}$. Besides, if $R^{\prime}(q^{{\dagger}})\geq\underline{w}/2$ it must hold $q^{{\dagger}}=q^{*}=1$.
Consequently, we can obtain that
Upper bounding PoA.  We next prove the upper bound of PoA.
By Lemma 5.5, we have that for any monopoly revenue $R^{*}\geq 0$, the largest first-best benchmark for one agent is achieved by the truncated equal-revenue distribution with monopoly revenue equal to $R^{*}$. We can obtain that
$U_{\textsc{fb}}=\int_{0}^{+\infty}1-F^{n}(z)\cdot{\mathrm{d}}z\leq R^{*}\cdot\left(1+\ln\frac{\bar{w}}{R^{*}}\right)\overset{(a)}{\leq}R^{*}\cdot(1+\ln\kappa),$
As a result, we can obtain that for any regular distribution,
Lower bounding PoDM and PoA.  We now analyze Example 5.5, which shows the lower bound part of the two ratios, i.e., $\mathrm{PoA}=\Omega_{\kappa}\left(\left(\log\kappa\right)^{2}\right)$ and $\mathrm{PoDM}=\Omega_{\kappa}\left(\log\kappa\right)$.
By definition, the virtual value for $z\in[z_{k},z_{k-1})$ is
$\phi(z)=R^{\prime}(q)=2^{K-k+1},$
Next, we consider the PoA. We have
$U_{\textsc{fb}}={}$ $\int_{0}^{z_{0}}1-F(z)\cdot{\mathrm{d}}z$
[…]
=$\overset{(a)}{=}{}$ $\sum_{k\in[2:K]}\frac{k-2+2^{1-k}}{K}\cdot\ln\left(2+\frac{1}{2^{k-1}-1}\right)$
$\overset{(b)}{\geq}{}$ $\sum_{2\in[K]}\frac{k-2}{K}\cdot\ln 2$
=$\overset{(c)}{=}{}$ $\frac{K^{2}-3K+3}{2K}\cdot\ln 2,$
@@ 5.2 Comparison under MHR Distributions
In this subsection, we bound the two ratios under identical expected rewards $r$ and MHR contribution distributions. Our first main result in Theorem 5.6 shows that for MHR contribution distributions, both ratios are constant—a significant improvement over the results obtained for regular distributions. Notably, unlike our findings for regular distributions in Theorem 7.1, Theorem 5.6 does not impose any restrictions on the support of the contribution distribution.
@@
For agents with identical expected rewards $r$ and MHR contribution distributions $F$, the tight PoDM is in $[{\mathrm{e}},3.033]$ and the tight PoA is ${\mathrm{e}}^{2}$. Moreover, if $n=1$, the tight PoDM is ${\mathrm{e}}$.
@@
Implications of Theorem 5.7 The upper bound in Theorem 5.7 depends on both the number of agents $n$ and the probability $F(0)$. Viewing the agents as a market of size $n$, the market quality can be measured by $1-F(0)$, the probability that an agent’s expected reward exceeds her cost. There are two interesting observations. Firstly, in a market of fixed quality, as the market size grows, the impact of the intermediary’s participation and the presence of agents’ private information on the utility of the principal diminishes to zero. Secondly, for two markets of the same size, this impact is smaller in the higher-quality market (higher $1-F(0)$) compared to lower-quality one.
@@
Proof overview of Theorem 5.6 Here, we outline the main ideas behind the proof of Theorem 5.6, which consists of three parts. In the first part, we start by scaling down the maximized utility $U^{*}$ to ensure that the revenue from anonymous pricing is at most ${\mathrm{e}}\cdot U^{*}$ (Lemma 5.10). Using the ratio between the first-best benchmark and the revenue from anonymous pricing (Lemma 5.11), we find that the upper bound of PoA is ${\mathrm{e}}^{2}$.
@@
In the second part, we apply another version of Lemma 5.10 to scale down $U^{*}$ and break the upper bound for PoDM into two multiplicative factors, one of which is provided by Lemma 5.12. For the other factor, we use its monotonicity for scaling. We calculate the upper bound for $n\leq 44$ , as shown in Table 3. For $n>44$, since both factors decrease, we conclude that the upper bound of PoDM is $3.033$. In the third part, we provide an example (Example 5.7) to illustrate the lower bounds of the two ratios. The lower bound part of Theorem 5.6 utilizes Example 5.7. The upper bound part utilizes Lemmas 5.8, 5.9, 5.10, 5.11 and 5.12.
@@
Proof overview of Theorem 5.7 Here we provide a high-level overview of the proof of Theorem 5.7. We first reference a conclusion that provides powerful tail-bounds with respect to the order statistics of the distribution (Lemma 5.13). Through this conclusion, we prove that the ratio of the first-best benchmark to the principal’s utility has an upper bound related to the contract and the expectation of the first-order statistic (Lemma 5.14). Then, we scale the principal’s optimal utility by selecting a specific contract, thereby obtaining an upper bound for PoA in relation to the number of agents $n$.
Example 5.7.
We know that $F_{r}$ is an MHR distribution with $F(\tilde{w})=0$ and $F(0)=1-1/K$.
Lemma 5.8 (Aggarwal et al., 2009, Lemma 1).
For any random variable $v$ that follows an MHR distribution $F$ with support in $[0,+\infty)$, $1-F(\theta)\geq\frac{1}{{\mathrm{e}}}$, where $\theta$ is the monopoly reserve price of the distribution $F$.
It is important to note that Lemma 5.8 holds only when the support of the distribution is non-negative. We next extend this conclusion to cases with arbitrary support sets.
Corollary 5.9.
Proof.
$1-F(\theta)={}$ ${\mathrm{e}}^{-\int_{-\infty}^{\theta}h(z)\cdot{\mathrm{d}}z}$
[…]
=$\overset{(b)}{=}{}$ $\frac{1}{{\mathrm{e}}}(1-F(0))$
[…]
A consequence of Corollary 5.9 is that the principal’s utility under a linear contract $\alpha$ can be lower-bounded as follows.
Lemma 5.10.
Proof.
where $z=r(1-\alpha)+\theta_{\alpha}$. Notice that by Lemma 5.2, the distribution $F^{k}$ is MHR for any $k\in\mathbb{N}$. We denote the monopoly reserve price of $F_{\alpha}^{k}$ as $\theta_{\alpha}^{(k)}$. Based on Corollary 5.9, we directly obtain that
$U(\alpha)={}$ $r(1-\alpha)(1-F_{\alpha}^{n}(\theta_{\alpha}))$
[…]
Moreover, when $k=n$, we have
$U(\alpha)\geq\frac{1}{{\mathrm{e}}}r(1-\alpha)(1-F^{n}(r(1-\alpha))).$
And when $k=1$, we have
Lemma 5.10 serves as a crucial tool for our subsequent analysis of the two ratios. Specifically, we will repeatedly utilize the expressions for $k=n$ and $k=1$ in our forthcoming proofs.
Lemma 5.11 (Dhangwatnotai et al., 2015, Theorem 3.11).
Lemma 5.12 (Jin et al., 2019a, Theorem 1).
Proof of Theorem 5.6..
We first show the upper bound parts utilizing the above lemmas and then show the lower bound part by analyzing Example 5.7.
$U^{*}\overset{(a)}{=}{}$ $\max_{\alpha\in[0,1]}U(\alpha)$
=$\overset{(c)}{=}{}$ $\frac{1}{{\mathrm{e}}}\theta^{(n)}(1-F^{n}(\theta^{(n)}))$
=$\overset{(e)}{=}{}$ $\frac{1}{{\mathrm{e}}^{2}}U_{\textsc{fb}},$
where equality (a) holds by definition of $U^{*}$, inequality (b) holds by Lemma 5.10, equality (c) holds be definition of $\theta^{(n)}$, inequality (d) holds by Lemma 5.11 and equality (e) holds by the definition of $U_{\textsc{fb}}$. Thus we have $\mathrm{PoA}=\frac{U_{\textsc{fb}}}{U^{*}}\leq{\mathrm{e}}^{2}$ by definition.
Upper bounding PoDM.  Next, we consider the upper bound of PoDM.
When, $n>1$, by Lemma 5.10, we have
It follows that
$\mathrm{PoDM}={}$ $\frac{U_{\textsc{sb}}}{U^{*}}$
[…]
Note that $Z_{n}(\cdot)$ is non-decreasing in $[0,1]$ and $1-F^{n}(\theta^{(n)})\geq\frac{1}{{\mathrm{e}}}$ by Lemma 5.8, we have
Lemma 5.12 provides the numerical results of $\tau_{n}$ for different $n$, which allows us to obtain the numerical results of the lower bound of PoDM. We offer the numerical results of the lower bound in Table 3 when $n\leq 44$. When $n\geq 44$, it can be verified that PoDM can never reach $3.033$ since $\tau_{n}$ and $Z_{n}((1-1/{\mathrm{e}})^{1/n}))$ are both decreasing as $n$ increases. Hence the lower bound of $3.033$ is reached when $n=7$.
$n$ 2 3 4 5 6 7 8 9 10 11
12 13 14 15 16 17 18 19 20 21 22
3.032 3.031 3.031 3.031 3.030 3.029 3.027 3.025 3.023 3.022 3.019
23 24 25 26 27 28 29 30 31 32 33
3.017 3.015 3.013 3.011 3.009 3.007 3.005 3.003 3.001 2.998 2.997
34 35 36 37 38 39 40 41 42 43 44
2.995 2.993 2.991 2.989 2.987 2.986 2.984 2.982 2.980 2.979 2.977
Table 3: The upper bounds of PoDM for MHR distributions when $n\leq 44$.
Lower bounding PoDM and PoA.  Now we analyze the lower bound part of the two ratios by considering Example 5.7. Formally, we show that for every fixed $n$ and any small real $\varepsilon>0$, there exists a parameter $K>1$ and a reward $r$ satisfying that $\mathrm{PoA}\geq{\mathrm{e}}^{2}-\varepsilon$ and $\mathrm{PoDM}\geq{\mathrm{e}}-\varepsilon$.
which does not rely on the parameter $K$.
Observe that for any $z\geq 0$, we have inequality
which gives the lower and upper bounds of $1-F_{r}^{n}(\cdot)$ as
$U^{*}={}$ $\max_{z\in[0,r]}\phi(z)(1-F_{r}^{n}(z))$
[…]
On the other hand,
$U_{\textsc{fb}}={}$ $\int_{0}^{r}(1-F^{n}(z))\cdot{\mathrm{d}}z$
[…]
and
$U_{\textsc{sb}}={}$ $\int_{0}^{r}\phi(z)(1-F_{r}^{n}(z))\cdot{\mathrm{d}}z$
[…]
Therefore, we can finally obtain the lower bounds
and
Finally, for an arbitrary large $r$, $\mathrm{PoA}\geq{\mathrm{e}}^{2}-\varepsilon$ and $\mathrm{PoDM}\geq{\mathrm{e}}-\varepsilon$. ∎
Lemma 5.13 (Giannakopoulos et al., 2021a, Lemma 3.3).
Lemma 5.14.
Proof.
By Lemma 5.10, we have:
Observe that the probability distribution function satisfies that
$\bar{F}(z)=\frac{F(z)-F(0)}{1-F(0)}\leq F(z)$
for every $z\in[0,\infty)$. Recall that $\bar{\mu}_{n}$ is the excepted value of the first-order statistic derived from $n$ i.i.d. random variables following the distribution $\bar{F}$. We have
Notice that $\bar{F}$ is MHR since $F$ is MHR. According to Lemma 5.13, we have
$\bar{F}(\delta\cdot\bar{\mu}_{n})\leq 1-{\mathrm{e}}^{-\delta\cdot\mathcal{H}_{n}},$
As a result, we obtain that
[…]
This completes the proof. ∎
Now we are ready to prove Theorem 5.7.
Proof of Theorem 5.7..
Consider the design of setting $\alpha^{{\dagger}}=1-\frac{\bar{\mu}_{n}}{r}(1-\frac{\ln\mathcal{H}_{n}}{\mathcal{H}_{n}})$. According to the definition of $\delta$ in Lemma 5.14, we have $\delta^{{\dagger}}=1-\frac{\ln\mathcal{H}_{n}}{\mathcal{H}_{n}}$. In addition, using the lower bound of $\frac{U(\alpha)}{U_{\textsc{fb}}}$ in Lemma 5.14, we obtain that
[…]
Therefore, since $\mathrm{PoDM}\leq\mathrm{PoA}$, we finally have
which completes the proof. ∎

### v2

@@ 4 Impact of Intermediary and Price of Double Marginalization
To this end, we introduce the concept of the price of double marginalization (PoDM), which captures the efficiency loss resulting from the intermediary’s presence. Specifically, PoDM quantifies the ratio between the principal’s utility under the second-best benchmark, in which she directly interacts with agents via an optimal DSIC mechanism, and the her utility obtained in our delegated mechanism design model with an intermediary.
@@
To evaluate the overall efficiency loss relative to the welfare-optimal outcome, we additionally consider the classical price of anarchy (PoA). This benchmark compares the principal’s utility under the first-best benchmark—which maximizes total welfare—to that in our model.
@@
where $U^{*}$ denotes the principal’s optimal utility in our model, and $U_{\textsc{fb}}$ denotes the first-best benchmark, corresponding to the maximum expected social welfare achievable when the principal can observe the realized cost profile and directly select the agent that maximizes welfare:
@@
The first-best and second-best benchmarks in the above definitions correspond, respectively, to the maximum welfare and the maximum revenue in a related single-item auction:
Since the contribution distributions $\bm{F}$ are regular by 2.3, we have the optimal mechanism in the second-best benchmark is the virtual welfare maximizer mechanism. Therefore, the second-best benchmark is equivalent to the expected revenue of the virtual welfare maximizer mechanism, i.e., $U_{\textsc{sb}}=\mathbb{E}_{\bm{w}\sim\bm{F}}\!\left[{\max\nolimits_{i\in[n]}\left\{{\left(\phi_{i}(w_{i})\right)^{+}}\right\}}\right]$.
Furthermore, the first-best benchmark is equivalent to the maximized utility of the principal in the scenario where the principal has direct access to agents and knows their cost realizations. Indeed, the assumption in this scenario implies that the principal can assign the task to the one with the lowest cost, thereby obtaining a utility equivalent to her contribution. Therefore, the first-best benchmark is equivalent to the expectation of the maximum value between 0 and the first order statistic $w_{\max}$ from the $n$ contribution, i.e., $U_{\textsc{fb}}=\mathbb{E}_{\bm{w}\sim\bm{F}}\!\left[{{\left(w_{\max}\right)^{+}}}\right]$.
@@ 4.1 Comparison under Regular Distributions
In this subsection, we bound both PoDM and PoA. Our main result in Theorem 4.1 shows that for a regular contribution distribution $F$, if the agents’ maximum possible contribution is at most $\kappa$ times their minimum possible contribution, then PoDM is tightly bounded by $\Theta(\log\kappa)$, while the PoA is tightly bounded by $\Theta((\log\kappa)^{2})$. [^15] We say a ratio is tight if, for all contribution distributions satisfying the given conditions and any number of agents $n$, the ratio does not exceed a certain value, while there exists at least one distribution and one $n$ where the ratio equals or approaches this value.
@@
Proof overview of Theorem 4.1. Here, we outline the main ideas behind the proof of Theorem 4.1, which consists of three parts. In the first part, we establish the upper bound for PoDM under two conditions with different arguments. For the case where $4\leq\bar{w}/\underline{w}\leq\kappa$, we apply the monotonicity of the virtual value to reduce the maximized utility $U^{*}$ (Lemma C.7). We then divide the quantile space of any contribution distribution into $2\left\lceil\log\kappa\right\rceil+1$ pieces. Next, using the monotonicity of the revenue curve for regular distributions, we find an upper bound for monopoly revenue $R^{*}$ and the ratio between the maximized utility $U^{*}$ and $R^{*}$. For the case where $1\leq\bar{w}/\underline{w}\leq 4$, we construct a quantile $q^{{\dagger}}$ representing the largest quantile where $R^{\prime}(q^{{\dagger}})\geq\underline{w}/2$. Then, by rescaling the integral form for $R^{*}$, we obtain a constant upper bound for the ratio between $U^{*}$ and $R^{*}$. We finally use the ratio of revenues from anonymous pricing and Myerson’s optimal auction (Lemma C.8) to establish an upper bound for the second-best benchmark $U_{\textsc{sb}}$. This process allows us to establish the upper bound for PoDM.
@@
In the second part, we define the truncated equal-revenue distribution (Definition C.3) and show that it achieves the maximum first-best benchmark for a single agent (Lemma C.9). Combining this with the upper bound for PoDM, we derive the upper bound for PoA.
@@
In the third part, we provide an example (Example C.2) to illustrate the lower bounds of the two ratios and demonstrate that the two ratios are tight. Specifically, the lower bound part of Theorem 4.1 utilizes Example C.2. The upper bound part utilizes Lemmas C.6, C.7, C.8 and C.9. The formal proof of Theorem 4.1 can be found in Section C.3.
@@ 4.2 Comparison under MHR Distributions
In this subsection, we bound both PoDM and PoA for MHR contribution distributions. Our first main result in Theorem 4.2 shows that for MHR contribution distributions, both ratios are constant—a significant improvement over the results obtained for regular distributions. Notably, unlike our findings for regular distributions in Theorem 4.1, Theorem 4.2 does not impose any restrictions on the support of the contribution distribution.
@@
For agents with identical expected rewards $r$ and MHR contribution distributions $F$, the tight PoDM is in ${\mathrm{e}},3.033]$ and the tight PoA is ${\mathrm{e}}^{2}$. Moreover, if $n=1$, the tight PoDM is $\mathrm{e}$}.
@@
Implications of Theorem 4.3. The upper bound in Theorem 4.3 depends on both the number of agents $n$ and the probability $F(0)$. Viewing the agents as a market of size $n$, the market quality can be measured by 1F(0)1-F(0), which is the probability that an agent’s expected reward exceeds her cost. There are two interesting observations. Firstly, in a market of fixed quality, as the market size grows, the impact of the intermediary’s participation and the presence of agents’ private information on the utility of the principal diminishes to zero. Secondly, for two markets of the same size, this impact is smaller in the higher-quality market (higher 1F(0)1-F(0)) compared to the lower-quality one.
@@
Proof overview of Theorem 4.2. Here, we outline the main ideas behind the proof of Theorem 4.2, which consists of three parts. In the first part, we start by scaling down the maximized utility $U^{*}$ to ensure that the revenue from anonymous pricing is at most ${\mathrm{e}}\cdot U^{*}$ (Lemma C.12). Using the ratio between the first-best benchmark and the revenue from anonymous pricing (Lemma C.13), we find that the upper bound of PoA is ${\mathrm{e}}^{2}$.
@@
In the second part, we apply another version of Lemma C.12 to scale down $U^{*}$ and break the upper bound for PoDM into two multiplicative factors, one of which is provided by Lemma C.14. For the other factor, we use its monotonicity for scaling. We calculate the upper bound for $n\leq 44$, as shown in Table 3. For $n>44$, since both factors decrease, we conclude that the upper bound of PoDM is $3.033$.
In the third part, we provide an example (Example C.4) to illustrate the lower bounds of the two ratios. The lower bound part of Theorem 4.2 utilizes Example C.4. The upper bound part utilizes Lemmas C.10, C.11, C.12, C.13 and C.14. The formal proof of Theorem 4.2 can be found in Section C.4
@@
Proof overview of Theorem 4.3. Here we provide a high-level overview of the proof of Theorem 4.3. We first reference a conclusion that provides powerful tail bounds with respect to the order statistics of the distribution (Lemma C.15). Through this conclusion, we prove that the ratio of the first-best benchmark to the principal’s utility has an upper bound related to the contract and the expectation of the first-order statistic (Lemma C.16). Then, we scale the principal’s optimal utility by selecting a specific contract, thereby obtaining an upper bound for PoA in relation to the number of agents $n$. The formal proof of Theorem 4.3 can be found in Section C.5